1)The thief grabbed the cell phone and ran at 6 meters per second straight down Ocean Street. When he was 120 meters down the street a police officer took off after him from the scene of the crime. The officer chased at the rate of 9 meters per second. How long will it take for the officer to catch the thief?

2)If $18,000 is invested at 5% for 6 years, find the future value if the interest is compounded daily.

3)The medical office purchased new computers at a cost of $1,855 at 14% add-on interest for 2 years. Find the APR (rounded to the nearest tenth of a percent) for the loan. (

4) Karen and Wayne need to buy a refrigerator because theirs just broke. Unfortunately, their savings account is depleted, and they will need to borrow money in order to buy a new one. Sears offers them an installment loan at 14% (add-on rate). If the refrigerator at Sears costs $1,500 plus 8% sales tax, and Karen and Wayne plan to pay for the refrigerator for 2 years, what is the monthly payment?

5)

The 9 members of the Board of Directors for the Eco-gro company voted to award a contract for new office equipment. The choices were (A) Anfernee, (D) Deebo, (L) Latifah, and (M) Moesha. Here are the preferences:

(ALMD)

(ADLM)

(LAMD)

(DALM)

3

2

4

3

Which company is selected by the Borda count method?

6)

The fraternity is electing a school president and there are four candidates: Kim (K), Selena (S), Carl (C), and Princess (P). The voter preferences are:

(CKPS)

(CKSP)

(PKSC)

(SKPC)

45

30

35

45

Who wins the election by the Hare method?

7)

The fraternity is electing a school president and there are four candidates: Kim (K), Selena (S), Carl (C), and Princess (P). The voter preferences are:

(CKPS)

(CKSP)

(PKSC)

(SKPC)

45

30

35

45

Does this violate the Condorcet Criterion and why?
Answer #1

1. The police officer will be able to catch the thief only if can travel the marginal distance between them ,which is here given as 120 meter.

Now as the thief and police officer both are going in them same direction therefore the relative speed of the police officer w.r.t the thief will be =speed of the police officer - speed of the thief=9-6=3m/s.

Hence time required to catch the thief will be

T= Marginal distance / Relative speed = 120/3=40sec.

Hence after 40 second the police officer may catch the thief.

2. as the interest is compounded daily we may call it as continuously compounded ,for continuously compounded interest we have amount

A=P*ert

here p=18000,r=.05,t=6.

therefore

A=18000*e.05*6=18000*1.349852=24297.458.

For the rest problem plse repost as at time we cant answer all owing to time constraint .